DETERMINE WHETHER (8,-9) IS A SOLUTION OF 8X+3Y=-1
Answer by richwmiller(9143)

Why are you yelling? Use both upper and lower case.
Just plug 8 in for x and -9 for y
8*8+3*-9=-1
64-27=-1
64=26
Obviously not equal so the point (8,-9) is not on the line 8x+3y=-1